OpenStudy (anonymous):

I have to find the solution that satisfies the initial condition y(0)=0 \[ln y + \frac{y^{2}}{2}= sin x +C\]

OpenStudy (anonymous):

does that mean I did the integration wrong? this is a continuation of the previous problem

OpenStudy (anonymous):

maybe

OpenStudy (anonymous):

\[\frac{dy}{dx}=\frac{ycosx}{1+y^{2}}\]

myininaya (myininaya):

So you did integration be use of separating the variables right?

OpenStudy (anonymous):

yes ma'am

OpenStudy (anonymous):

\[\int \frac{1+y^{2}}{y}dy= sinx +C\]

OpenStudy (anonymous):

\[ln y + \frac{y^{2}}{2}= sin x +C\]

myininaya (myininaya):

hmm... y(0)=0 will not work as the initial condition

OpenStudy (anonymous):

sorry y(0)= 1

myininaya (myininaya):

oh! :) replace x with 0 and y with 1 :) that will work

myininaya (myininaya):

And then solve for C

OpenStudy (anonymous):

yeah!!!! .5?

myininaya (myininaya):

yep yep that is the value for C :)

myininaya (myininaya):

So just write your equation over replacing C with .5

OpenStudy (anonymous):

sounds good, thanks!